A truck and a car moving with the same kinetic energy are brought to rest by application of brakes which provide equal retarding Force for both. The truck has 8 tyres and weights 2 times more than the cars which has 4 tyres. What can you say about the distance in which the two vehicles stop?

Grade:10

2 Answers

jitender
6 years ago
Let kinetic energies of trick and car be Kt and Kc respectively.Now , Kt = Kc (given) ____*1W=♤K......(work done = change in kinetic energy...( because potential energy is same throughout the journey ( ♤h= 0)))(♤= change = delta)Also W = F•dLet truck travel x distance and car y aftér application of brakes. ♤Kt = F•x _____2*♤Kc = F•y _____3*From 1* , 2* , and 3*Fx = Fyx = y.....Both travel same distance on application of Same regarding force if they were moving with same initial kinetic energies.
Athul sekhar
6 years ago
bothkinetic energy of the truck = Kinetic energy of car(given)
Velocity of both the vehicles are same (given) v1=v2
Let time both the vehicles be t1 and t2 
1/2v1 t1^2 = 1/2v2 v2^2 
By cancelling v1 , v2 and ½ , ½. 
We get
t1^2=t2^2
t1=t2
From the above statement , we can say that
both the vehicle will stop at the same time